# Thermally driven two-sphere microswimmer with internal feedback control

### v1

@@ VI Example
So far, we have not yet specified the functional forms of the spring potential $U(x)$ and the friction coefficients $\gamma_{i}(x)$. As an explicit example, we choose the harmonic potential energy $U(x)=Kx^{2}/2$, where $K>0$ is the spring constant, and assume that the friction coefficients have the form $\gamma_{1}(x)=\gamma_{2}(x)=\gamma[1+\delta\tanh(x/w)]$, where $\gamma$ is a constant friction coefficient, $\delta$ is a dimensionless parameter satisfying $|\delta|<1$ (recall $\gamma_{i}(x)>0$), and $w$ is a length characterizing the change of the friction coefficients Kumar08 ; Baule08 . With this choice, the steady-state distribution function in Eq. (13) and the average velocity in Eq. (15) become
@@
respectively, where $\varepsilon=Kw^{2}/(k_{\mathrm{B}}T_{1})$ is a dimensionless parameter. Note that Eq. (21) is simply a Gaussian distribution function. Since the dimensionless integral in Eq. (22) is positive (which will be evaluated numerically below), $\langle V\rangle$ is finite only when $\delta\neq 0$ and $\theta\neq 1$. Moreover, the temperature ratio $\theta$ determines the direction of locomotion, i.e., $\langle V\rangle<0$ when $\delta>0$ and $\theta>1$, or vice versa. Similarly, by calculating the average heat flow $\langle\dot{Q}_{1}\rangle$ in Eq. (18), we obtain the average entropy production rate in Eq. (20) as
@@
in which the integral is dimensionless.
In Fig. 2(a), we numerically plot the dimensionless average velocity $\langle\hat{V}\rangle=\langle V\rangle\gamma/\sqrt{Kk_{\mathrm{B}}T_{1}}$ in Eq. (22) as a function of $\delta$ for different values of $\theta=T_{2}/T_{1}>1$ when $\varepsilon=1$. To confirm our analytical prediction, we have also performed numerical simulations of the coupled stochastic equations in Eqs. (4) and (5). These Langevin equations were discretized according to the Itô convention. We find a good agreement between the theoretical and numerical results. Notice that $\langle\hat{V}\rangle$ vanishes when either $\theta=1$ or $\delta=0$, as predicted by Eq. (22). Moreover, $\langle\hat{V}\rangle$ is an odd function of $\delta$ because the sign of $\delta$ determines the direction of locomotion. In Fig. 2(b), on the other hand, we numerically plot the dimensionless average entropy production rate $\langle\hat{\dot{\sigma}}\rangle=\langle\dot{\sigma}\rangle\gamma T_{1}/(Kw)^{2}$ in Eq. (23) as a function of $\delta$. For $\theta>1$, $\langle\dot{\sigma}\rangle$ is positive and is an even function of $\delta$. As a result, the entropy production rate is minimized at $\delta=0$ and the corresponding expression is given by $\langle\dot{\sigma}\rangle_{\rm min}=k_{\mathrm{B}}K(\theta-1)^{2}/(2\gamma\theta)\geq 0$.
At this stage, it is worth evaluating the efficiency of the microswimmer defined by Golestanian08 ; Sou21
@@
where $T^{\ast}=(T_{1}+T_{2})/2$ is the average temperature although either $T_{1}$ or $T_{2}$ can also be used. The factor $2\gamma$ in the numerator reflects the fact the microswimmer consists of two spheres with the same average friction coefficient $\gamma$. In Fig. 3, we color-plot the efficiency $\eta$ as a function of $\delta$ and $\theta$. The efficiency $\eta$ vanishes when $\delta=0$ and is an even function of $\delta$ when $\theta$ is fixed. It is worth noting that, in the equilibrium limit of $\theta\rightarrow 1$, the efficiency $\eta$ is finite because both $\langle V\rangle$ and $\langle\dot{\sigma}\rangle$ vanish in this limit [see Eqs. (22) and (23)].
VII Discussion and summary
In our model, the state-dependent friction coefficients play an essential role. The change in the friction coefficient can be caused, for example, by the change in the particle size. In this sense, the current model has a similarity to the model of “pushmepullyou” by Avron et al. Avron05 or the two-sphere model by Pandey et al. Pandey12 . The main difference in our model is that the spheres have different temperatures, and thermal fluctuations cause the locomotion in a stochastic manner. The internal heat flow combined with the feedback control through the position-dependent friction coefficient is the main driving force for maintaining persistent locomotion. In general, there is a contribution to the total entropy production rate coming from the center-of-mass diffusion, which is neglected in Eq. (24Sou21 . In the above efficiency, we have considered the entropy production rate due to the heat flow [see Eq. (16)] that causes the unidirectional motion.
@@ VII Discussion and summary
In summary, we have studied the locomotion of a thermally driven elastic two-sphere microswimmer that has position-dependent friction coefficients. Such a microswimmer can acquire nonzero average velocity in the steady state due to the heat flow between the spheres. We have obtained the entropy production rate and further estimated the efficiency of the microswimmer. The proposed self-propulsion mechanism emphasizes the importance of active matter and biological systems driven by information. In the future, we will calculate the time-correlation functions of the microswimmer to quantitatively discuss the degree of broken time-reversal symmetry Li24 .
@@ Acknowledgements.
We thank Z. Xiong for the useful discussion. Z.H.  L.H.  and S.K. acknowledge the support by the National Natural Science Foundation of China (Nos. 12104453, 22273067, 12274098, and 12250710127). Y.H. acknowledges support from JSPS Overseas Research Fellowships (Grant No. 202460086). S.K. acknowledges the startup grant of Wenzhou Institute, University of Chinese Academy of Sciences (No. WIUCASQD2021041). K.Y and S.K. acknowledge the support by the Japan Society for the Promotion of Science (JSPS) Core-to-Core Program “Advanced core-to-core network for the physics of self-organizing active matter” (No. JPJSCCA20230002). J.L. and Z.Z. contributed equally to this work.

### v2

@@ VI Example
So far, we have not yet specified the functional forms of the spring potential $U(x)$ and the friction coefficients $\gamma_{i}(x)$. As an example, we choose the harmonic potential energy $U(x)=Kx^{2}/2$, where $K>0$ is the spring constant, and assume that the friction coefficients have the form Kumar08 ; Baule08
$\gamma_{1}(x)=\gamma_{2}(x)=\gamma[1+\delta\tanh(x/w)],$ (22)
where $\gamma$ is a constant friction coefficient, $\delta$ is a dimensionless feedback strength parameter satisfying $|\delta|<1$ (recall $\gamma_{i}(x)>0$), and $w$ is a length characterizing the change of the friction coefficients. In this case, the steady-state distribution function in Eq. (14) and the average velocity in Eq. (16) become
@@
respectively, where $\varepsilon=Kw^{2}/(k_{\mathrm{B}}T_{1})$ is the dimensionless energy parameter. Note that Eq. (23) is simply a Gaussian distribution function. Since the dimensionless integral in Eq. (24) is positive (which will be evaluated numerically below), $\langle V\rangle$ is non-zero only when $\delta\neq 0$ and $\theta\neq 1$. Moreover, the temperature ratio $\theta$ determines the direction of locomotion, i.e., $\langle V\rangle<0$ when $\delta>0$ and $\theta>1$, or vice versa. Similarly, by calculating the average heat flow $\langle\dot{Q}_{1}\rangle$ in Eq. (19), we obtain the average entropy production rate in Eq. (21) as
@@
which includes another dimensionless integral.
In Fig. 2(a), we numerically plot the dimensionless average velocity $\langle\hat{V}\rangle=\langle V\rangle\gamma/\sqrt{Kk_{\mathrm{B}}T_{1}}$ from Eq. (24) as a function of the feedback strength parameter $\delta$ for different values of the temperature ratio $\theta=T_{2}/T_{1}>1$ when $\varepsilon=1$. To confirm our analytical prediction, we have also performed numerical simulations of the coupled stochastic equations in Eqs. (5) and (6). These Langevin equations were discretized according to the Itô convention. We find a good agreement between the theoretical and numerical results. Notice that $\langle\hat{V}\rangle$ vanishes when either $\theta=1$ or $\delta=0$, as shown in Eq. (24). Moreover, $\langle\hat{V}\rangle$ is an odd function of $\delta$ because the sign of $\delta$ determines the direction of locomotion. In Fig. 2(b), on the other hand, we numerically plot the dimensionless average entropy production rate $\langle\hat{\dot{\sigma}}\rangle=\langle\dot{\sigma}\rangle\gamma T_{1}/(Kw)^{2}$ from Eq. (25) as a function of the feedback strength $\delta$. For $\theta>1$, $\langle\dot{\sigma}\rangle$ is positive and is an even function of $\delta$. As a result, the entropy production rate is minimized at $\delta=0$ and the corresponding expression is given by $\langle\dot{\sigma}\rangle_{\rm min}=k_{\mathrm{B}}K(\theta-1)^{2}/(2\gamma\theta)\geq 0$.
At this stage, it is worth evaluating the efficiency of the current microswimmer defined by Golestanian08 ; Sou21
@@
where $T^{\ast}=(T_{1}+T_{2})/2$ is the average temperature although either $T_{1}$ or $T_{2}$ can also be used. The factor $2\gamma$ in the numerator reflects the fact that the microswimmer consists of two spheres with the same average friction coefficient $\gamma$. In Fig. 3, we color-plot the efficiency $\eta$ as a function of the feedback strength parameter $\delta$ and the temperature ratio $\theta=T_{2}/T_{1}$. The efficiency $\eta$ vanishes when $\delta=0$ and is an even function of $\delta$ when $\theta$ is fixed. It is worth noting that, in the equilibrium limit of $\theta\rightarrow 1$, the efficiency $\eta$ is finite because both $\langle V\rangle$ and $\langle\dot{\sigma}\rangle$ vanish in this limit [see Eqs. (24) and (25)].
In general, there is a contribution to the total entropy production rate coming from the center-of-mass diffusion Sou21 , which is neglected in Eq. (26). Here, we have only considered the entropy production rate due to the heat flow [see Eq. (17)] that causes the unidirectional locomotion.
VII Summary and Discussion
@@
To summarize, we have studied the locomotion of a thermally driven elastic two-sphere microswimmer that has position-dependent friction coefficients. Such a microswimmer can acquire non-zero average velocity in the steady state due to the heat flow between the spheres. We have obtained the entropy production rate and further estimated the efficiency of the microswimmer. The proposed self-propulsion mechanism emphasizes the importance of active matter and biological systems driven by internal feedback control. In the future, we will calculate the time-correlation functions of the microswimmer to quantitatively discuss the degree of broken time-reversal symmetry Li24 .
In our model, the state-dependent friction coefficients play an essential role. The change in the friction coefficient can be caused, for example, by the change in the particle size. In this sense, the current model has a similarity to the model of “pushmepullyou” by Avron et al. Avron05 or the two-sphere model by Pandey et al. Pandey12 . The main difference in our model is that the spheres have different temperatures and thermal fluctuations cause the locomotion in a stochastic manner. The internal heat flow combined with the feedback control through the position-dependent friction coefficient is the main driving force for persistent locomotion.
The other important assumption of the model is that the statistical properties of the two spheres are characterized by two different temperatures. Such a micromachine is conceptually motivated by Feynman’s Brownian ratchet Feynmanbook or a model system that interacts with two thermal environments SekimotoBook . More recently, a similar non-equilibrium dimer model confined between two walls was discussed to calculate probability flux loops that demonstrate the broken detailed balance when the temperatures are different Battle16 ; Gnesotto18 ; Li19 . In the previous thermally driven three-sphere microswimmer model Hosaka17 , all the spheres were assumed to have different temperatures.
Although it is technically challenging to introduce different temperatures in a moving microswimmer, one possible way is to use a chemically heterogeneous particle, such as a Janus particle, under laser irradiation, which induces inhomogeneous temperature distribution on the particle Li19 . By further introducing the internal deformation degree of freedom Yang14 , we expect that the proposed thermally driven two-sphere microswimmer can be experimentally realized. It should be mentioned, however, that any difference in the fluctuation (both thermal and non-thermal) between the two spheres is sufficient to induce a non-zero velocity of the two-sphere microswimmer Kumar08 ; Baule08 . Hence, rather than the thermodynamic temperature itself, an effective temperature that reflects the microswimmer’s internal fluctuation is more important for its persistent locomotion.
@@ Acknowledgements.
We thank Z. Xiong for the useful discussions. Z.H., L.H.  and S.K. acknowledge the support by the National Natural Science Foundation of China (Nos. 12104453, 22273067, 12274098, and 12250710127). Y.H. acknowledges support from JSPS Overseas Research Fellowships (Grant No. 202460086). S.K. acknowledges the startup grant of Wenzhou Institute, University of Chinese Academy of Sciences (No. WIUCASQD2021041). K.Y and S.K. acknowledge the support by the Japan Society for the Promotion of Science (JSPS) Core-to-Core Program “Advanced core-to-core network for the physics of self-organizing active matter” (No. JPJSCCA20230002). J.L. and Z.Z. contributed equally to this work.